OpenStudy (dj3strella):

Question 2.2. Find the slope and y-intercept of each line. y = x – 3

OpenStudy (anonymous):

Do the same thing I told you. This is the basic formula: y = mx + b m = slope slope = 1 b = -3

OpenStudy (dj3strella):

so when x is by its self its 1x?

OpenStudy (anonymous):

Yeah, we don't ever write 1x, just x. When it's just x it means it really is 1x. Is that helpful?

OpenStudy (dj3strella):

yes thank you @pabloecortez07

OpenStudy (anonymous):

No problem, ask me another question, glad to help.

OpenStudy (anonymous):

This one's a bit more complicated, but same concept. The basic formula is : y = mx + b So let's try to get it in that form. We can subtract 4x on both sides, that gives us: 6y = -4x + 12 We can divide by 6 on both sides to get y by itself. y = -4/6x + 2 We are almost done, but we can still simplify -4/6, since that's really -2/3 So our final answer is: y = -2/3x + 2 Did that help?

OpenStudy (dj3strella):

yes

OpenStudy (anonymous):

Awesome